How would I solve a question like this: 6.21 g of iron(III) sulfate was dissolved in a solution of 300 mL. What is the concentration, in mol/L, of the sulfate ion in solution?
1
Expert's answer
2020-12-21T03:48:47-0500

We can solve this using the formula; n=CV

But we need to find the number of moles first Using the formula; mole=mass/molar mass

Molar mass of Iron(iii)sulfate=400g/mol

Mole=6.21/400=0.016moles

Coverting 300mL to litres we have 0.3L

If n=CV ,then C=n/V

C=0.016/0.3

C=0.053M